OpenStudy (anonymous):

Evaluate each expression for x = -2 and y = 3 11. 6xy + 2x2 + 4x 12. (3x + y2)2

OpenStudy (anonymous):

6(-2)(3)+2(-2)2+4(-2) solve

OpenStudy (anonymous):

(3(-2)+(3)2)2 solve

OpenStudy (anonymous):

128 for the first one

OpenStudy (anonymous):

make that -128

OpenStudy (anonymous):

the 2nd is 0 because2*0=0

OpenStudy (anonymous):

thanks @jazzy688

OpenStudy (anonymous):

your welcome itry to be the best

OpenStudy (anonymous):

How did you get -128 for the first one? I got -52. Could you explain? Thanks!

OpenStudy (anonymous):

I meant can you post the math as to how you arrived at your answer? Here's how I got -52, maybe you can show me where I went wrong. 6 * (-2) * 3 = -36 + 2 * (-2) * 2 = -8 + 4 * (-2) = -8 -36 + -8 + -8 = -52 Thanks

OpenStudy (anonymous):

 we both are wrong miss calculations

OpenStudy (anonymous):

I think you left out a 2 here: (-2)2+4(-2) as it is 2* the Variable X * 2, not 2 x 2?